# On the growth of actions of free products

### v1

@@ Abstract.
If $G$ is a finitely generated group and $X$ a $G$-set, the growth of the action of $G$ on $X$ is the function that measures the largest cardinality of a ball of radius $n$ in the Schreier graph $\Gamma(G,X)$. In this note we consider the following stability problem: if $G,H$ are finitely generated groups admitting a faithful action of growth bounded above by a function $f$, does the free product $G\ast H$ also admit a faithful action of growth bounded above by $f$? We show that the answer is positive under additional assumptions, and negative in general. In the negative direction, our counter-examples are obtained with $G$ either the commutator subgroup of the topological full group of a minimal and expansive homeomorphism of the Cantor space; or $G$ a Houghton group. In both cases, the group $G$ admits a faithful action of linear growth, and we show that $G\ast H$ admits no faithful action of subquadratic growth provided $H$ is non-trivial. In the positive direction, we describe a class of groups that admit actions of linear growth and is closed under free products and exhibit examples within this class, among which the Grigorchuk group.
@@ 1. Introduction
It is not hard to see that the class $\mathcal{C}(f)$ is stable under taking finitely generated subgroups, and under passing to a finite index overgroup [LBMB22a, §1.3]. We use the special notation $\mathcal{C}(\mathrm{lin})$ for the distinguished case $f(n)\simeq n$, which is the slowest possible growth for a faithful action of an infinite group. The class $\mathcal{C}(\mathrm{lin})$ is richer than it might look at first sight. It is not hard to check that it contains all virtually abelian groups. The fact that non-abelian free groups belong to $\mathcal{C}(\mathrm{lin})$ is well-known and can be traced back to Schreier [Sch27]. A famous example of a group in $\mathcal{C}(\mathrm{lin})$ is the Grigorchuk group (and all Grigorchuk groups $G_{\omega}$ from [Gri84]), see Bartholdi and Grigorchuk [BG00]. Many other interesting groups of dynamical origin are naturally given by an action of linear growth (and hence belong to $\mathcal{C}(\mathrm{lin})$): topological full groups of homeomorphisms of the Cantor set, and Nekrashevych’s fragmentation of dihedral groups [Nek18]. More examples of groups in $\mathcal{C}(\mathrm{lin})$ include the lamplighter group $C_{p}\wr\mathbb{Z}$, the Houghton groups, or Neumann’s groups from [Neu37].
@@
In [Sal21] the third author investigated the class of subgroups of the topological full group of full shifts over $\mathbb{Z}$ (which is contained in $\mathcal{C}(\mathrm{lin})$), and gave sufficient conditions under which graph products of such groups remain inside that class [Sal21, Theorem 3]. That result implies in particular that the class $\mathcal{C}(\mathrm{lin})$ contains all right angled Artin groups [Sal21, Theorem 1], and hence all their subgroups as well. Combined with results of many authors, this is a vast source of examples of groups in $\mathcal{C}(\mathrm{lin})$, as many groups are known to embed (upon to passing to a finite index subgroup) in a right angled Artin groups. For example, this is the case for instance surface groups, and more generally any hyperbolic group acting geometrically on a CAT(0) cube complex, as follows from Agol’s final step in the solution of the virtual Haken conjecture [Ago13] combined with the previous results of Haglund and Wise [HW08].
@@
Motivated by this last class of examples, we consider here the question whether the class $\mathcal{C}(\mathrm{lin})$ (or the classes $\mathcal{C}(f)$ for more general $f$) is closed under free products, or more generally graph products. There is a natural strategy to attempt to answer that question affirmatively. Given two groups $G_{1}$ and $G_{2}$ and $X_{i}$ a $G_{i}$-set, $i=1,2$, one may consider any set $X=X_{1}\sqcup_{A_{1}\cong_{\phi}A_{2}}X_{2}$ obtained by identifying suitable subsets $A_{1}\subset X_{1}$ and $A_{2}\subset X_{2}$ via a bijection $\phi\colon A_{1}\to A_{2}$. The action of each $G_{i}$ extends to $X$ trivially outside of the copy of $X_{i}$ in $X$. This defines an action of the free product $G=G_{1}\ast G_{2}$. On the one hand, if the subsets along which the $X_{i}$ are glued are chosen wisely, one can hope to control the growth function $\operatorname{vol}_{G,X}(n)$ in terms of $\operatorname{vol}_{G_{i},X_{i}}(n),i=1,2$. On the other hand, if the gluing is sufficiently generic, the action of $G$ on $X$ is likely to be faithful. The question therefore becomes whether, for $G_{1},G_{2}$ in $\mathcal{C}(f)$, the tension between these two conditions can be conciliated. A naive implementation of this strategy (suitably adapted to more general graph products) gives the following. We refer to §2.1 for the definition of graph products.
Proposition 1 (Proposition 2.7).
@@
Let $G_{1},\ldots,G_{k}\in\mathcal{C}(f)$. Then the free product $G=G_{1}\ast\cdots\ast G_{k}$ belongs to $\mathcal{C}(g)$, where $g(n)=nf(n)$. More generally, any graph product of the $G_{i}$’s belongs to $\mathcal{C}(g)$.
@@
Proposition 1 is very often not optimal. In many cases, appropriate choices allow to show that a graph product of groups in $\mathcal{C}(f)$ remains in $\mathcal{C}(f)$. In §2.3 we discuss a sufficient condition for this to be the case. That condition is an elaboration of a condition considered in [Sal21]. Proposition 2.13 provides in particular a direct and elementary proof of the fact that right-angled Artin groups belong to $\mathcal{C}(\mathrm{lin})$ (which essentially follows the same lines as the proof in [Sal21], but avoids the technicalities coming from the full group setting). More generally, Proposition 2.13 allows to exhibit examples of graph products in the class $\mathcal{C}(\mathrm{lin})$:
Proposition 2.
@@
The free product $G_{1}\ast\cdots\ast G_{k}$ belongs to $\mathcal{C}(\mathrm{lin})$ whenever each free factor $G_{i}$ is isomorphic to one of the following:
@@ 1.1. The main results
The main goal of this note is to provide examples that show that the bound obtained from the simple construction from Proposition 1 is sharp in general, even for free products. We consider two families of examples.
The first are topological full groups of minimal group actions on the Cantor set, more precisely their alternating subgroups in the sense of Nekrashevych [Nek19] (see §3.4 for definitions). Following [LBMB22a], we shall say that a finitely generated group $G$ has a Schreier growth gap $f(n)$ if every faithful $G$-set $X$ satisfies $\operatorname{vol}_{G,X}(n)\succcurlyeq f(n)$.
Theorem 1.
@@
A relevant special case in the previous theorem is $\mathcal{G}=\mathbb{Z}$ (in that case $G$ coincides with the commutator subgroup of the topological full group, by the results of Matui [Mat06]). In that case we obtain examples of groups $G\in\mathcal{C}(\mathrm{lin})$ such that $G\ast H$ has a Schreier growth gap $n^{2}$ for any non-trivial group $H$. The quadratic gap is optimal in general, as Proposition 1 says that $G\ast H$ has a faithful action of growth $\simeq n^{2}$ for instance when $H$ is cyclic and non-trivial.
Our second family of examples are the Houghton groups $H_{r},r\geq 2$. The group $H_{r}$ is finitely generated and belongs to $\mathcal{C}(\mathrm{lin})$ for every $r\geq 2$ (see §3.3).
Theorem 2.
@@
1.2. Outline of the proof of Theorems 1 and 2
@@
. These two results share the same proof mechanism. It relies on the description of the confined subgroups of the groups $G$ in their statements, obtained in previous works of the first two authors. Recall that a subgroup $H$ of a group $G$ is confined if the set of conjugates of $H$ does not accumulate on the trivial subgroup in the space $\operatorname{Sub}(G)$ of subgroups of $G$, endowed with the Chabauty topology. When $G$ is finitely generated, convergence in the space $\operatorname{Sub}(G)$ can be interpreted in the space of marked Schreier graphs on the corresponding coset spaces $G/H$. The study of confined subgroups of a group $G$ leads (among other applications) to results on the growth function $\operatorname{vol}_{G,X}(n)$ of $G$-actions, such as the existence of a Schreier growth gap [MB18, LBMB23, LBMB22a]. For $G$ as in Theorem 1 or 2, a classification of the confined subgroups of $G$ has been obtained respectively in [MB18] and [LBMB22a]. Using these results, we show that every faithful action of $G$ whose growth is close to the growth of the natural defining action of $G$ (respectively on the Cantor set or on the bouquet of rays), must be “almost” conjugate to it. A common feature of the two situations is that the group $G$ contains elements whose support is a very sparse subset of the Schreier graph of the natural action. We exploit this fact and the previous result about actions of small growth of $G$ to show that in the graph for any faithful action of $G\ast H$, there must be short jumps between regions that are far in the graph of the restricted action of $G$. This is that phenomenon that forces an additional factor $n$ in the growth. The detailed proofs are given in Section 3.
@@
It is worth comparing the case of the Grigorchuk group in Proposition 2 with Theorem 1. The Grigorchuk group and topological full groups share various common features: in particular they appear through a micro-supported action on a compact space, a condition which plays an important role in the study of confined subgroups (indeed the confined subgroups of the Grigorchuk group are also understood [LBMB23]). Also Theorem 1 is applicable to many fragmentations of dihedral groups in the sense of Nekrashevych [Nek18] (a family of groups to which the Grigorchuk group also belongs). Despite these similarities, here they exhibit an opposite behaviour. The main reason for this difference is the fact that its action on the vertices of the binary tree satisfies a condition that we call large displacement (Definition 2.9), which means that every non-trivial element of the group must move some vertex of the tree by a distance (on the Schreier graph of the corresponding level) comparable to the total diameter of the Schreier graph (Proposition 2.14). Although the proof of this fact is not difficult, it relies on rather specific features of the Grigorchuk group.
@@ 2.1. Graph products
Let $I$ be a set, and let $\Delta$ denote the diagonal in $I^{2}$. Let $c:I^{2}\setminus\Delta\to\left\{0,1\right\}$ such that $c(i,j)=c(j,i)$ for all $i\neq j$. If $(G_{i})_{i\in I}$ a family of groups indexed by $I$, the graph product $P=(G_{i})_{i\in I}^{c}$ of the family $(G_{i})_{i\in I}$ associated to $c$ is the quotient of the free product $\ast_{I}G_{i}$ by the relations $[G_{i},G_{j}]$ for all $i,j$ such that $c(i,j)=1$. We say that $P$ is a finite graph product if $I$ is finite.
@@ 2.2. Preliminaries
Let $q\geq 1$. Let $(X_{1},\ldots,X_{q})=(X_{k})_{k\leq q}$ be a $q$-tuple of graphs, and for every $k\leq q$ let $(e_{k},s_{k})$ be two distinct points of $X_{k}$. The gluing of the $q$-tuple $(X_{k},e_{k},s_{k})_{k\leq q}$ is the graph obtained by taking the disjoint union $X_{1}\sqcup\cdots\sqcup X_{q}$, and identifying $s_{k}$ and $e_{k+1}$ for every $11\leq k\leq q-$ 1. It is denoted $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$.
@@
Every ball $B$ of radius $n$ in $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ intersects at most $12n+$ 1 members of $(X_{k})_{k\leq q}$, and is covered by $12n+$ 1 balls of radius at most $n$ within these $X_{k}$. The statement follows. ∎
@@
Now suppose that $(G_{i})_{i\in I}$ is a family of groups, and for every $i\in I$ there is an action of $G_{i}$ on $X$ for every $X\in\mathcal{F}_{i}$. Let $P=(G_{i})_{i\in I}^{c}$ be the graph product of $(G_{i})_{I}$ associated to $c$. For every $c$-admissible $(X_{k},e_{k},s_{k})_{k\leq q}$, there is a natural $G_{i}$-action on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ that extends the $G_{i}$-action on each $X_{k}$ such that $i(X_{k})=i$, that is defined by declaring that $G_{i}$ acts trivially on $X_{k}\setminus\left\{e_{k},s_{k}\right\}$ for every $k$ such that $i(X_{k})\neq i$. This is well defined because of the property that no two consecutive $X_{k}$ can be in $\mathcal{F}_{i}$, guaranteed by the assumption $i(X_{k})\neq i(X_{k+1})$ in Definition 2.3. This defines an action of the free product $\ast_{I}G_{i}$ on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$. The fact that $(X_{k},e_{k},s_{k})_{k\leq q}$ is $c$-admissible ensures that if $c(i,j)=1$, then the groups $G_{i}$ and $G_{j}$ act on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ with disjoint support. Hence the action of $\ast_{I}G_{i}$ on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ factors through the graph product $P=(G_{i})_{i\in I}^{c}$.
@@
Suppose for every $i\in I$ the group $G_{i}$ is generated by a subset $S_{i}$, and that the graph structure on each $X\in\mathcal{F}_{i}$ is the Schreier graph of the $G_{i}$-action on $X$. Then the graph $\mathcal{G}((X_{k},e_{k},s_{k}))$ is essentially the Schreier graph of the action of $P=(G_{i})_{i\in I}^{c}$ associated to the generating subset $\cup_{I}S_{i}$. The only difference is that loops should be added at the places where the action of the $G_{i}$’s has been extended in a trivial way.
@@
Suppose that for every $i\in I$, the $G_{i}$-action on $\bigsqcup_{X\in\mathcal{F}_{i}}X$ is faithful. Then the action of the graph product $P=(G_{i})_{i\in I}^{c}$ on $\mathcal{G}(\mathcal{F},c)$ is faithful.
@@
where for each $s$ we have that $g_{s}$ is non-trivial and there is $i_{s}\in I$ (necessarily unique) such that $g_{s}\in G_{i_{s}}$. Among all such decomposition, we choose one such that $n$ is minimal. Since $g$ is non-trivial, $n\geq 1$. We define a sequence $(r_{1},\ldots,r_{q})$ inductively by setting $r_{1}=1$, and defining $r_{k+1}$ as the smallest $s\geq r_{k}+1$ such that $c(i(s),i(r_{k}))=0$ (equivalently, $G_{i(s)}$ does not commute with $G_{i(r_{k})}$). We also define $r_{q+1}=n+1$. Minimality of $n$ implies that for every $1\leq k\leq q$ and every $r_{k}+1\leq s\leq r_{k+1}-1$, we have $g_{s}\notin G_{i(r_{k})}$. Let $\gamma_{k}=g_{r_{k+1}-1}\cdots g_{r_{k}}$ for $1\leq k\leq q$, so that $g=\gamma_{q}\cdots\gamma_{1}$. Since $G_{i}$ acts faithfully on $\bigsqcup_{X\in\mathcal{F}_{i}}X$ for every $i\in I$, for every $k\leq q$ one can find $X_{k}\in\mathcal{F}_{i_{r_{k}}}$ and $x_{k}\in X_{k}$ such that $g_{r_{k}}(x_{k})\neq x_{k}$. We set $e_{k}=x_{k}$, $s_{k}=g_{r_{k}}(x_{k})$. We look at how the element $g$ acts on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$. Since $g_{s}$ does not belong to $G_{i(r_{k})}$ for every $r_{k}+1\leq s\leq r_{k+1}-1$, $g_{s}$ acts trivially on $s_{k}$. Hence $\gamma_{k}(e_{k})=s_{k}$ for every $k\leq q$, and hence it follows that $g(e_{1})=s_{q}$. So if $q=1$ then $g$ acts non-trivially since $e_{1}\neq s_{1}$. If $q\geq 2$ then $X_{q}$ and $X_{1}$ are either disjoint (when $q\geq 3$), or intersect only along $s_{1}=e_{2}$ if $q=2$. Hence in all cases the element $g$ moves $e_{1}$. So for every non-trivial element $g$ of $P$, there exists a $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ on which $g$ acts non-trivially. So $P$ acts faithfully on $\mathcal{G}(\mathcal{F},c)$. ∎
Recall that $\mathcal{C}(f)$ has been defined in the introduction (Definition 1.1).
@@
Let $f\colon\mathbb{N}\to\mathbb{N}$, and let $(G_{i})_{I}$ be a finite collection of groups in the class $\mathcal{C}(f)$. Then any graph product $P=(G_{i})_{i\in I}^{c}$ belongs to $\mathcal{C}(g)$ with $g(n)=nf(n)$.
@@
For each $i\in I$ there is a $G_{i}$-set $X_{i}$ such that the $G_{i}$-action on $X_{i}$ is faithful and verifies $\operatorname{vol}_{G_{i},X_{i}}(n)\preccurlyeq f(n)$. The statement then follows from Lemma 2.2 and Lemma 2.6 (and Remark 2.5), applied to $\mathcal{F}_{i}=\left\{X_{i}\right\}$ and $\mathcal{F}=\sqcup_{I}\mathcal{F}_{I}$. ∎

### v2

@@ Abstract.
If $G$ is a finitely generated group and $X$ a $G$-set, the growth of the action of $G$ on $X$ is the function that measures the largest cardinality of a ball of radius $n$ in the (possibly non-connected) Schreier graph $\Gamma(G,X)$. We consider the following stability problem: if $G,H$ are finitely generated groups admitting a faithful action of growth bounded above by a function $f$, does the free product $G\ast H$ also admit a faithful action of growth bounded above by $f$? We show that the answer is positive under additional assumptions, and negative in general. In the negative direction, our counter-examples are obtained with $G$ either the commutator subgroup of the topological full group of a minimal and expansive homeomorphism of the Cantor space, or $G$ a Houghton group. In both cases, the group $G$ admits a faithful action of linear growth, and we show that $G\ast H$ admits no faithful action of subquadratic growth provided $H$ is non-trivial. In the positive direction, we describe a class of groups that admit actions of linear growth and is closed under free products and exhibit examples within this class, among which the Grigorchuk group.
@@
Dedicated with admiration to Slava Grigorchuk on the occasion of his 70th birthday.
@@ 1. Introduction
It is not hard to see that the class $\mathcal{C}(f)$ is stable under taking finitely generated subgroups, and under passing to a finite index overgroup [LBMB22b, §1.3]. We use the special notation $\mathcal{C}(\mathrm{lin})$ for the distinguished case $f(n)\simeq n$, which is the slowest possible growth for a faithful action of an infinite group. The class $\mathcal{C}(\mathrm{lin})$ is richer than it might look at first sight. It is not hard to check that it contains all virtually abelian groups. The fact that non-abelian free groups belong to $\mathcal{C}(\mathrm{lin})$ is well-known and can be traced back to Schreier [Sch27]. A famous example of a group in $\mathcal{C}(\mathrm{lin})$ is the Grigorchuk group (and all Grigorchuk groups $G_{\omega}$ from [Gri84]), see Bartholdi and Grigorchuk [BG00]. Many other interesting groups of dynamical origin are naturally given by an action of linear growth (and hence belong to $\mathcal{C}(\mathrm{lin})$): topological full groups of homeomorphisms of the Cantor set, and Nekrashevych’s fragmentation of dihedral groups [Nek18]. More examples of groups in $\mathcal{C}(\mathrm{lin})$ include the lamplighter group $C_{p}\wr\mathbb{Z}$, the Houghton groups, or Neumann’s groups from [Neu37] (the latter are defined by an action whose orbits are all finite, yet having linear growth).
@@
In [Sal21] the third author investigated the class of subgroups of the topological full group of full shifts over $\mathbb{Z}$ (which is contained in $\mathcal{C}(\mathrm{lin})$), and gave sufficient conditions under which graph products of such groups remain inside that class [Sal21, Theorem 3]. That result implies in particular that the class $\mathcal{C}(\mathrm{lin})$ contains all right angled Artin groups [Sal21, Theorem 1], and hence all their subgroups as well. Combined with results of many authors, this is a vast source of examples of groups in $\mathcal{C}(\mathrm{lin})$, as many groups are known to embed (upon to passing to a finite index subgroup) in some right angled Artin group. For example, this is the case for instance surface groups, and more generally any hyperbolic group acting geometrically on a CAT(0) cube complex, as follows from Agol’s final step in the solution of the virtual Haken conjecture [Ago13] combined with the previous results of Haglund and Wise [HW08].
@@
Motivated by this last class of examples, we consider here the question whether the class $\mathcal{C}(\mathrm{lin})$ (or the classes $\mathcal{C}(f)$ for more general $f$) is closed under free products, or more generally graph products. There is a natural strategy to attempt to answer that question affirmatively. Given two groups $G_{1}$ and $G_{2}$ and $X_{i}$ a $G_{i}$-set, $i=1,2$, one may consider any set $X=X_{1}\sqcup_{A_{1}\cong_{\phi}A_{2}}X_{2}$ obtained by identifying suitable subsets $A_{1}\subset X_{1}$ and $A_{2}\subset X_{2}$ via a bijection $\phi\colon A_{1}\to A_{2}$. The action of each $G_{i}$ extends to $X$ as a trivial action outside of the copy of $X_{i}$ in $X$. This defines an action of the free product $G=G_{1}\ast G_{2}$. On the one hand, if the subsets along which the $X_{i}$ are glued are chosen wisely, one can hope to control the growth function $\operatorname{vol}_{G,X}(n)$ in terms of $\operatorname{vol}_{G_{i},X_{i}}(n),i=1,2$. On the other hand, if the gluing is sufficiently generic, the action of $G$ on $X$ is likely to be faithful. The question therefore becomes whether, for $G_{1},G_{2}$ in $\mathcal{C}(f)$, the tension between these two conditions can be conciliated. A naive implementation of this strategy (suitably adapted to more general graph products) gives the following. We refer to §2.1 for the definition of graph products.
Proposition 1.2 (Proposition 2.7).
@@
Let $k\geq 1$ and $G_{1},\ldots,G_{k}\in\mathcal{C}(f)$. Then the free product $G=G_{1}\ast\cdots\ast G_{k}$ belongs to $\mathcal{C}(g)$, where $g(n)=nf(n)$. More generally, any graph product of the $G_{i}$’s belongs to $\mathcal{C}(g)$.
@@
Proposition 1.2 is very often not optimal. In many cases, appropriate choices allow to show that a graph product of groups in $\mathcal{C}(f)$ remains in $\mathcal{C}(f)$. In §2.3 we discuss a sufficient condition for this to be the case. That condition is an elaboration of a condition considered in [Sal21]. Proposition 2.13 provides in particular a direct and elementary proof of the fact that right-angled Artin groups belong to $\mathcal{C}(\mathrm{lin})$ (which essentially follows the same lines as the proof in [Sal21], but avoids the technicalities coming from the full group setting). More generally, Proposition 2.13 allows to exhibit examples of graph products in the class $\mathcal{C}(\mathrm{lin})$:
Proposition 1.3.
@@
For $k\geq 1$, the free product $G_{1}\ast\cdots\ast G_{k}$ belongs to $\mathcal{C}(\mathrm{lin})$ whenever each free factor $G_{i}$ is isomorphic to one of the following:
@@ 1.1. The main results
The main goal of this paper is to provide examples that show that the bound obtained from the simple construction from Proposition 1.2 is sharp in general, even for free products. We consider two families of examples.
The first are topological full groups of minimal group actions on the Cantor set, more precisely their alternating subgroups in the sense of Nekrashevych [Nek19] (see §3.4 for definitions). Following [LBMB22b], we shall say that a finitely generated group $G$ has a Schreier growth gap $f(n)$ if every faithful $G$-set $X$ satisfies $\operatorname{vol}_{G,X}(n)\succcurlyeq f(n)$.
Theorem 1.4.
@@
A relevant special case in the previous theorem is $\mathcal{G}=\mathbb{Z}$ (in that case $G$ coincides with the commutator subgroup of the topological full group, by the results of Matui [Mat06]). In that case we obtain examples of groups $G\in\mathcal{C}(\mathrm{lin})$ such that $G\ast H$ has a Schreier growth gap $n^{2}$ for any non-trivial group $H$. The quadratic gap is optimal in general, as Proposition 1.2 says that $G\ast H$ has a faithful action of growth $\simeq n^{2}$ for instance when $H$ is cyclic and non-trivial.
Our second family of examples is the family of Houghton groups $H_{r},r\geq 2$. Recall that the group $H_{2}$ is defined as the group of permutations of $\mathbb{Z}$ that coincide with a translation outside a finite set; see §3.3 for the definition of $H_{r}$. Each group $H_{r}$ is finitely generated and belongs to $\mathcal{C}(\mathrm{lin})$.
Theorem 1.5.
@@
1.2. Outline of the proof of Theorems 1.4 and 1.5
@@
These two results share the same proof mechanism. It relies on the description of the confined subgroups of the groups $G$ in their statements, obtained in previous works of the first two authors. Recall that a subgroup $H$ of a group $G$ is confined if the set of conjugates of $H$ does not accumulate on the trivial subgroup in the space $\operatorname{Sub}(G)$ of subgroups of $G$, endowed with the Chabauty topology. When $G$ is finitely generated, convergence in the space $\operatorname{Sub}(G)$ can be interpreted in the space of marked Schreier graphs on the corresponding coset spaces $G/H$. The study of confined subgroups of a group $G$ leads (among other applications) to results on the growth function $\operatorname{vol}_{G,X}(n)$ of $G$-actions, such as the existence of a Schreier growth gap [MB18, LBMB23, LBMB22b]. For $G$ as in Theorem 1.4 or 1.5, a classification of the confined subgroups of $G$ has been obtained respectively in [MB18] and [LBMB22b]. Using these results, we show that every faithful action of $G$ whose growth is close to the growth of the natural defining action of $G$ (respectively on the Cantor set or on the bouquet of rays), must be “almost” conjugate to it. A common feature of the two situations is that the group $G$ contains elements whose support is a very sparse subset of the Schreier graph of the natural action. We exploit this fact and the previous result about actions of small growth of $G$ to show that in the graph for any faithful action of $G\ast H$, there must be short jumps between regions that are far in the graph of the restricted action of $G$. This is that phenomenon that forces an additional factor $n$ in the growth. The detailed proofs are given in Section 3.
@@
It is worth comparing the case of the Grigorchuk group in Proposition 1.3 with Theorem 1.4. The Grigorchuk group and topological full groups share various common features: in particular they appear through a micro-supported action on a compact space, a condition which plays an important role in the study of confined subgroups (indeed the confined subgroups of the Grigorchuk group are also understood [LBMB23]). Also Theorem 1.4 is applicable to many fragmentations of dihedral groups in the sense of Nekrashevych [Nek18] (a family of groups to which the Grigorchuk group also belongs). Despite these similarities, here they exhibit an opposite behaviour. The main reason for this difference is the fact that the action of the Grigorchuk group on the vertices of the binary tree satisfies a condition that we call orbits of controlled diameter (Definition 2.9), which means that every non-trivial element of the group must move some vertex of the tree by a distance comparable to the total diameter of the Schreier graph of the action on the corresponding level of the tree (Proposition 2.14). Although the proof of this fact is not difficult, it relies on rather specific features of the Grigorchuk group.
Acknowledgements
We thank the referees for their careful reading of the paper.
@@ 2.1. Graph products
Let $I$ be a set, and let $\Delta$ denote the diagonal in $I^{2}$. Let $c$ be a map $I^{2}\setminus\Delta\to\left\{0,1\right\}$ such that $c(i,j)=c(j,i)$ for all $i\neq j$. Note the data of $c$ is equivalent to the data of an undirected graph with vertex set $I$ with no loop and simple edges. If $(G_{i})_{i\in I}$ a family of groups indexed by $I$, the graph product $P=\mathsf{GP}((G_{i})_{I};c)$ of the family $(G_{i})_{i\in I}$ associated to $c$ is the quotient of the free product $\Asterisk_{I}G_{i}$ by the relations $G_{i},G_{j}]$ for all $i,j$ such that $c(i,j)=1$. We say that $P$ is a finite graph product if $I$ is finite.
@@ 2.2. Preliminaries
Let $q\geq 1$. Let $(X_{1},\ldots,X_{q})=(X_{k})_{k\leq q}$ be a $q$-tuple of graphs, and for every $k\leq q$ let $(e_{k},s_{k})$ be two distinct points of $X_{k}$. The gluing of the $q$-tuple $(X_{k},e_{k},s_{k})_{k\leq q}$ is the graph obtained by taking the disjoint union $X_{1}\sqcup\cdots\sqcup X_{q}$, and identifying $s_{k}$ and $e_{k+1}$ for every $1\leq k\leq q-1$. It is denoted $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$.
@@
Every ball $B$ of radius $n$ in $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ intersects at most $2n+1$ members of $(X_{k})_{k\leq q}$, and is covered by $2n+1$ balls of radius at most $n$ within these $X_{k}$. The statement follows. ∎
@@
Now suppose that $(G_{i})_{i\in I}$ is a family of groups, and for every $i\in I$ there is an action of $G_{i}$ on $X$ for every $X\in\mathcal{F}_{i}$. Let $P=\mathsf{GP}((G_{i})_{I};c)$ be the graph product of $(G_{i})_{I}$ associated to $c$. For every $c$-admissible $(X_{k},e_{k},s_{k})_{k\leq q}$, there is a natural $G_{i}$-action on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ that extends the $G_{i}$-action on each $X_{k}$ such that $i(X_{k})=i$, that is defined by declaring that $G_{i}$ acts trivially on $X_{k}\setminus\left\{e_{k},s_{k}\right\}$ for every $k$ such that $i(X_{k})\neq i$. This is well defined because of the property that no two consecutive $X_{k}$ can be in $\mathcal{F}_{i}$, guaranteed by the assumption $i(X_{k})\neq i(X_{k+1})$ in Definition 2.3. This defines an action of the free product $\Asterisk_{I}G_{i}$ on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$. The fact that $(X_{k},e_{k},s_{k})_{k\leq q}$ is $c$-admissible ensures that if $c(i,j)=1$, then the groups $G_{i}$ and $G_{j}$ act on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ with disjoint support. Hence the action of $\Asterisk_{I}G_{i}$ on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ factors through the graph product $P=\mathsf{GP}((G_{i})_{I};c)$.
@@
Suppose for every $i\in I$ the group $G_{i}$ is generated by a subset $S_{i}$, and that the graph structure on each $X\in\mathcal{F}_{i}$ is the Schreier graph of the $G_{i}$-action on $X$. Then the graph $\mathcal{G}((X_{k},e_{k},s_{k}))$ is essentially the Schreier graph of the action of $P=\mathsf{GP}((G_{i})_{I};c)$ associated to the generating subset $\bigcup_{I}S_{i}$. The only difference is that loops should be added at the places where the action of the $G_{i}$’s has been extended in a trivial way.
@@
Suppose that for every $i\in I$, the $G_{i}$-action on $\bigsqcup_{X\in\mathcal{F}_{i}}X$ is faithful. Then the action of the graph product $P=\mathsf{GP}((G_{i})_{I};c)$ on $\mathcal{G}(\mathcal{F},c)$ is faithful.
@@
where for each $s$ we have that $g_{s}$ is non-trivial and there is a (necessarily unique) $i_{s}\in I$ such that $g_{s}\in G_{i_{s}}$. Among all such decompositions, we choose one such that $n$ is minimal. Since $g$ is non-trivial, $n\geq 1$. We define a sequence $(r_{1},\ldots,r_{q})$ inductively by setting $r_{1}=1$, and defining $r_{k+1}$ as the smallest $s\geq r_{k}+1$ such that $c(i(s),i(r_{k}))=0$ (equivalently, $G_{i(s)}$ does not commute with $G_{i(r_{k})}$). We also define $r_{q+1}=n+1$. Minimality of $n$ implies that for every $1\leq k\leq q$ and every $r_{k}+1\leq s\leq r_{k+1}-1$, we have $g_{s}\notin G_{i(r_{k})}$. Let $\gamma_{k}=g_{r_{k+1}-1}\cdots g_{r_{k}}$ for $1\leq k\leq q$, so that $g=\gamma_{q}\cdots\gamma_{1}$.
Since $G_{i}$ acts faithfully on $\bigsqcup_{X\in\mathcal{F}_{i}}X$ for every $i\in I$, for every $k\leq q$ one can find $X_{k}\in\mathcal{F}_{i_{r_{k}}}$ and $x_{k}\in X_{k}$ such that $g_{r_{k}}(x_{k})\neq x_{k}$. We set $e_{k}=x_{k}$, $s_{k}=g_{r_{k}}(x_{k})$. We look at how the element $g$ acts on $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$. Since $g_{s}$ does not belong to $G_{i(r_{k})}$ for every $r_{k}+1\leq s\leq r_{k+1}-1$, $g_{s}$ acts trivially on $s_{k}$. Hence $\gamma_{k}(e_{k})=s_{k}$ for every $k\leq q$, and hence it follows that $g(e_{1})=s_{q}$. So if $q=1$ then $g$ acts non-trivially since $e_{1}\neq s_{1}$. If $q\geq 2$ then $X_{q}$ and $X_{1}$ are either disjoint (when $q\geq 3$), or intersect only along $s_{1}=e_{2}$ if $q=2$. Hence in all cases the element $g$ moves $e_{1}$. So for every non-trivial element $g$ of $P$, there exists a $\mathcal{G}((X_{k},e_{k},s_{k})_{k\leq q})$ on which $g$ acts non-trivially. So $P$ acts faithfully on $\mathcal{G}(\mathcal{F},c)$. ∎
Recall that the class $\mathcal{C}(f)$ has been defined in the introduction (Definition 1.1).
@@
Let $f\colon\mathbb{N}\to\mathbb{N}$, and let $(G_{i})_{I}$ be a finite collection of groups in the class $\mathcal{C}(f)$. Then any graph product $P=\mathsf{GP}((G_{i})_{I};c)$ belongs to $\mathcal{C}(g)$ with $g(n)=nf(n)$.
@@
For each $i\in I$ there is a $G_{i}$-set $X_{i}$ such that the $G_{i}$-action on $X_{i}$ is faithful and verifies $\operatorname{vol}_{G_{i},X_{i}}(n)\preccurlyeq f(n)$. The statement then follows from Lemma 2.2 and Lemma 2.6 (and Remark 2.5), applied to $\mathcal{F}_{i}=\left\{X_{i}\right\}$ and $\mathcal{F}=\bigsqcup_{I}\mathcal{F}_{I}$. ∎